# Governance of Ledger-Anchored Decentralized Identifiers

### v1

@@ I Introduction
According to the core principles of SSI [5], any authoritative action related to a human-centered digital identity should be carried out by the very same individual. For example, a DID subject binding a new public key to their DID. But in practice and in particular with device IDs, a DID often needs to be controlled by others, even if the DID refers to a human being. For example, parents wish to control their minor children’s DID or the head of human resources wishes to control the DIDs of the company’s employees [6]. To enable this, authority over a DID can be shared with or delegated to others. This article explores the different concepts and realizations for governing authority over a ledger-anchored DID. It investigates how a DID subject appoints deputies, what control rights a deputy possesses over a DID, and how potentially multiple deputies coordinate control of a commonly governed DID among each other. It also presents an exemplary DID-conform identifier management system realized using Smart Contracts (SCs) on Ethereum. The prototype permits to appoint multiple deputies with different authorizations for a DID, and its governance of DIDs is transparent and fully customizable via SCs and extensible and interchangeable during operation.
@@ II Decentralized Identifier
did:iota IOTA n/a DID is fully controlled by a single ”state controller”. ”Governors” can revoke a DDO and appoint and dismiss state controllers. The DID subject may coincide with the state controller.
@@ III Related Work
Most DID methods do not require a declaration of DID controllers in the DDO. Instead, they rely on VDR-technology-specific mechanisms to manage authorizations and to check whether an entity trying to update a DDO in the VDR is authorized to do so. It is common to grant only a single entity DDO update rights. Prominent examples are did:ethr[14], did:bnb[15], did:indy[16], and did:sov[17]. However, this is a rather inflexible model as it limits the application of DIDs to domains where a digital identity is only controlled by a single entity. Only a few DID methods support multiple DID controllers, e.g., did:algo[18], did:sol[19], and did:ebsi[20], but each DID controller is granted full DDO update rights. This does not follow the principle of least privilege as in certain cases a deputy is assigned more DDO update power than actually needed. A DID subject, on the other side, might trust deputies to update their own verification methods in the DID subject’s DDO after a key rotation but not to appoint additional deputies that the DID subject might not know. More fine-granular authorizations can be defined in did:factom[21] by giving each DID controller a priority. Such a DID controller is only permitted to update verification methods of a DID controller with the same or lower priority. In did:iota[22], an entity with the role state controller is permitted to fully modify the DDO. The right to replace a state controller by another remains only with the entity possessing the governor role.
@@ IV Governance
At first, governance is broadly defined as the management and enforcement of control. This is inspired by DuPont’s suggestion to interpret it in the realm of decentralized autonomous organizations (DAOs) [24] as ”…a mechanism that sets institutional rules and incentives, or the strategic exercise of power…[25]. Yet, it remains challenging to provide a more suitable or tailored definition without using a specific lens [26] or without focusing on a particular aspect of the system under investigation. This article examines the governance of control over a ledger-anchored DID at the ledger’s application layer. Control over such a DID refers to the authority to update its ledger-anchored DDO. To gain control of a DID, an entity must be appointed as a DID controller by an already authorized entity, such as another DID controller. The latter, therefore, plays a key role in delegating control and determining who else is authorized to control the DID and to update, besides others, the governance policies in the DDO. Consequently, whoever has update power over the ledger-anchored DDO, governs the DID.
@@
The lifecycle of a ledger-anchored DDO is described here with four states: CREATED, ANCHORED, UPDATED, and INVALIDATED. The creation deals with the generation of at least one asymmetric key pair, the DID and DDO. It happens ideally locally at the DID subject’s premises, off-chain. The creation can also be taken care of by an entity different to the DID subject that acts as the DID subject’s deputy in the digital realm. This step is followed by anchoring the DID in the VDR respectively persisting the DDO in a distributed ledger so that the DID is linked to a single, unique, tamper-proof, and verifiable version of the DDO. If anything changes regarding the DID, such as keys are rotated for security purposes, the DDO is updated in the ledger to ensure that others become immediately aware of the latest version. If a DID subject ceases to exist and the DID is not needed anymore then the DDO is invalidated by deleting it from a redactable ledger or by flagging the DID as invalid. The lifecycle is illustrated in Figure 2.
@@ IV-A Policy Definition
After all, it lies in the hands of the ledger-based DID method to prescribe how a governance policy has to look like in the DDO. A ledger-based DID method is also permitted to completely forgo governance policies in the DDO and instead leave it to the ledger’s access control and its policies to decide about DDO updates. This, however, requires the ledger’s access control layer to become aware of the concept of DIDs. But enlisting governance policies in the DDO and enforcing them through ledger- or VDR-specific mechanisms offers advantages such as transparency, interoperability, and auditability. By publishing governance policies in a standardized form, interoperability is facilitated, as different ledgers can consistently interpret and enforce these in a general applicable form. It lays the foundation for DID governance to become assured and transferable across different ledgers, ledger technologies, and non-ledger VDR solutions. In addition, explicit update authorizations provide a clear history of permissions, which is important for auditability and regulatory compliance.
@@ IV-C3 Programmable Coordination
With group control, the DID specification permits entities to coordinate a DDO update among each other and submit it finally under the roof of a common and enlisted DID controller. If the coordination happens off-chain, resulting in a potentially multi-party-signed DDO update, then it lies in the hands of the group members to arrange their coordination securely and trustfully using techniques such as Multi-party Computation[35], Threshold Signature Schemes[36], or similar. Coordination can also be conducted transparently through the ledger, similar to the collaborative administration model for access control policies proposed in [37]. With programmable and decentrally executable logic in the form of SCs, all voting schemes [38] used in on-chain governance models for DAOs could serve as potential coordination mechanisms among accredited entities. An entity authorized to do so would announce a DDO update proposal transparently via an SC. Entities with dedicated voting rights for the DDO are then capable to approve or reject a DDO update proposal via the same SC. Examples of potential voting mechanism are weighted, n-out-of-m, unanimous and delegated voting, with or without time restrictions, anonymously or non-anonymously, and with either a policy-based authorization proof to vote or by means of tokens or VCs.
@@ IV-C4 Group-specific Coordination
A coordination mechanism may apply globally to the DID method or, if namable in the DDO, only in the scope of a single DID. If DID controllers of a single DID are dividable into groups by tagging each enlisted DID controller with the group name, then it becomes possible to apply different group-specific coordination mechanisms for a single DID. For example, a group of DID controllers with full update authority might need to agree unanimously on a critical governance-related DDO update while it is sufficient for another group with the authority to add an alternative DID subject identifier to reach an n-out-of-m consensus on the DDO update.
@@ VI Conclusion and Future Work
Allowing multiple entities to commonly control a ledger-anchored DID is a powerful feature. It extends the range of applications for DIDs to include use cases in which complex custodial, ownership, legal, and administrative relationships among natural persons, legal entities, and non-human entities need to be technically enforceable through the governance of the associated DID-based digital identities. The discussion about today’s deployed update governance models and on the framework set by the current DID specification, and the exploration of the possible enhancements and extensions of the framework highlight the crucial role governance policies in the DDO play for the flexibility and applicability of a DID method and its interoperability with other DID methods. Equally important is how these policies are technically enforced by the underlying VDR to ensure transparency, auditability, and as a consequence the trustworthiness of a ledger-based realization of a DID method. The proof of concept based on SCs demonstrates, as a first of its kind, the technical feasibility to realize a DID-conform and decentralized identifier management system where a DID is governed by multiple entities and where - at the same time - the governance is fully customizable and transparently replaceable during the DID’s lifetime.

### v2

@@ I Introduction
According to the core principles of SSI [5], any authoritative action related to a human-centered digital identity should be carried out by the very same individual. For example, a DID subject binding a new public key to their DID. But in practice and in particular with device IDs, a DID often needs to be controlled by others, even if the DID refers to a human being. For example, parents wish to control their minor children’s DID, or the head of human resources wishes to control the DIDs of the company’s employees [6]. To enable this, authority over a DID can be shared with or delegated to others. This article explores the different concepts and realizations for governing authority over a ledger-anchored DID. It investigates how a DID subject appoints deputies, what control rights a deputy possesses over a DID, and how potentially multiple deputies coordinate control of a commonly governed DID among each other. It also presents an exemplary DID-conform identifier management system realized using Smart Contracts (SCs) on Ethereum. The prototype permits to appoint multiple deputies with different authorizations for a DID, and its governance of DIDs is transparent and fully customizable via SCs and extensible and interchangeable during operation.
@@ II Decentralized Identifier
did:iota IOTA DID can be fully controlled by multiple entities. One controller makes an update proposal, others approve it, and the update is executed as soon as a predefined threshold is reached.
@@ III Related Work
Most DID methods do not require a declaration of DID controllers in the DDO. Instead, they rely on VDR-technology-specific mechanisms to manage authorizations and to check whether an entity trying to update a DDO in the VDR is authorized to do so. It is common to grant only a single entity DDO update rights. Prominent examples are did:ethr[14], did:bnb[15], did:indy[16], and did:sov[17]. However, this is a rather inflexible model as it limits the application of DIDs to domains where a digital identity is only controlled by a single entity. Only a few DID methods support multiple DID controllers, e.g., did:algo[18], did:sol[19], and did:ebsi[20], but each DID controller is granted full DDO update rights. This does not follow the principle of least privilege as in certain cases a deputy is assigned more DDO update power than actually needed. A DID subject, on the other hand, might trust deputies to update their own verification methods in the DID subject’s DDO after a key rotation but not to appoint additional deputies that the DID subject might not know. More fine-granular authorizations can be defined in did:factom[21] by giving each DID controller a priority. Such a DID controller is only permitted to update verification methods of a DID controller with the same or lower priority. In did:iota[22], a controller is able to make an update proposal. Other controllers of the same DDO approve the proposal. It is accepted and the update is executed as soon as a voting threshold is reached.
@@ IV Governance
Governance can be broadly defined as the management and enforcement of control. This is inspired by DuPont’s suggestion to interpret it in the realm of decentralized autonomous organizations (DAOs) [24] as ”…a mechanism that sets institutional rules and incentives, or the strategic exercise of power…[25]. Yet, it remains challenging to provide a more suitable or tailored definition without using a specific lens [26] or without focusing on a particular aspect of the system under investigation. This article examines the governance of control over a ledger-anchored DID at the ledger’s application layer. Control over such a DID refers to the authority to update its ledger-anchored DDO. To gain control of a DID, an entity must be appointed as a DID controller by an already authorized entity, such as another DID controller. The latter, therefore, plays a key role in delegating control and determining who else is authorized to control the DID and to update, besides others, the governance policies in the DDO. Consequently, whoever has update power over the ledger-anchored DDO, governs the DID.
@@
The lifecycle of a ledger-anchored DDO is described here with four states: CREATED, ANCHORED, UPDATED, and INVALIDATED. The creation deals with the generation of at least one asymmetric key pair, the DID and DDO. It happens ideally locally at the DID subject’s premises, off-chain. The creation can also be taken care of by an entity different from the DID subject that acts as the DID subject’s deputy in the digital realm. This step is followed by anchoring the DID in the VDR respectively persisting the DDO in a distributed ledger so that the DID is linked to a single, unique, tamper-proof, and verifiable version of the DDO. If anything changes regarding the DID, such as keys are rotated for security purposes, the DDO is updated in the ledger to ensure that others become immediately aware of the latest version. If a DID subject ceases to exist and the DID is not needed anymore then the DDO is invalidated by deleting it from a redactable ledger or by flagging the DID as invalid. The lifecycle is illustrated in Figure 2.
@@ IV-A Policy Definition
After all, it lies in the hands of the ledger-based DID method to prescribe how a governance policy must look like in the DDO. A ledger-based DID method is also permitted to completely forgo governance policies in the DDO and instead leave it to the ledger’s access control and its policies to decide about DDO updates. This, however, requires the ledger’s access control layer to become aware of the concept of DIDs. But enlisting governance policies in the DDO and enforcing them through ledger- or VDR-specific mechanisms offers advantages such as transparency, interoperability, and auditability. By publishing governance policies in a standardized form, interoperability is facilitated, as different ledgers can consistently interpret and enforce these in a general applicable form. It lays the foundation for DID governance to become assured and transferable across different ledgers, ledger technologies, and non-ledger VDR solutions. In addition, explicit update authorizations provide a clear history of permissions, which is important for auditability and regulatory compliance.
@@ IV-C3 Programmable Coordination
With group control, the DID specification permits entities to coordinate a DDO update among each other and submit it finally under the roof of a common and enlisted DID controller. If the coordination happens off-chain, resulting in a potentially multi-party-signed DDO update, then it lies in the hands of the group members to arrange their coordination securely and trustfully using techniques such as Multi-party Computation[35], Threshold Signature Schemes[36], or similar. Coordination can also be conducted transparently through the ledger, like the collaborative administration model for access control policies proposed in [37]. With programmable and decentrally executable logic in the form of SCs, all voting schemes [38] used in on-chain governance models for DAOs could serve as potential coordination mechanisms among accredited entities. An entity authorized to do so would announce a DDO update proposal transparently via an SC. Entities with dedicated voting rights for the DDO are then capable to approve or reject a DDO update proposal via the same SC. Examples of potential voting mechanism are weighted, n-out-of-m, unanimous and delegated voting, with or without time restrictions, anonymously or non-anonymously, and with either a policy-based authorization proof to vote or by means of tokens or VCs.
@@ IV-C4 Group-specific Coordination
A coordination mechanism may apply globally to the DID method or, if namable in the DDO, only in the scope of a single DID. If the DID controllers of a single DID can be divided into groups by tagging each registered controller with a group name, it becomes possible to apply group-specific coordination mechanisms for that DID. For example, a group of DID controllers with full update authority might need to agree unanimously on a critical governance-related DDO update while it is sufficient for another group with the authority to add an alternative DID subject identifier to reach an n-out-of-m consensus on the DDO update.
@@ VI Conclusion and Future Work
Allowing multiple entities to commonly control a ledger-anchored DID is a powerful feature. It extends the range of applications for DIDs to include use cases in which complex custodial, ownership, legal, and administrative relationships among natural persons, legal entities, and non-human entities need to be technically enforceable through the governance of the associated DID-based digital identities. The discussion about today’s deployed update governance models and on the framework set by the current DID specification, and the exploration of the possible enhancements and extensions of the framework highlight the crucial role governance policies in the DDO play for the flexibility and applicability of a DID method and its interoperability with other DID methods. Equally important is how these policies are technically enforced by the underlying VDR to ensure transparency, auditability, and consequently the trustworthiness of a ledger-based realization of a DID method. The proof of concept based on SCs demonstrates, as a first of its kind, the technical feasibility to realize a DID-conform and decentralized identifier management system where a DID is governed by multiple entities and where - at the same time - the governance is fully customizable and transparently replaceable during the DID’s lifetime.